# Rational motives on pro-algebraic stacks

### v3

@@ 2.1 Pullback formalisms
${F(Y_{j})}$${F_{K}(Y)}$${\lim_{K}F_{*}\circ Y}$${F(X_{j})}$${F_{K}(X)}$${\lim_{K}F_{*}\circ X}$$\scriptstyle{\textup{ins}_{Y}^{j}}$$\scriptstyle{f_{j}^{*}}$$\scriptstyle{f^{*}}$$\scriptstyle{h_{Y}}$$\scriptstyle{f^{*}}$$\scriptstyle{\textup{ins}_{X}^{j}}$$\scriptstyle{h_{X}}$
for any $j\in K$. First, we show that $p_{i}^{Y}\circ f^{*}$ commutes with colimits. Indeed, it is enough to see that $p_{i}^{Y}\circ f^{*}\circ h_{X}\circ\textup{ins}_{X}^{j}$ commutes with colimits, then is equivalent to the induced map $F_{K}(X)\rightarrow F(Y_{i})$ in $\Pr^{L}$, which necessarily commutes with colimits. By Lemma 2.20 (2), we have
$p_{i}^{X}\circ f^{*}\circ h_{Y}\circ\textup{ins}_{Y}^{j}\simeq p_{i}^{Y}\circ h_{X}\circ\textup{ins}_{X}^{j}\circ f_{j}^{*}\simeq\mathop{\textup{colim}}_{i\rightarrow k,j\rightarrow k}X_{ik*}\circ X_{jk}^{*}\circ f_{j}^{*}.$
By construction each of the functors $x_{ik*},x_{jk}^{*},f_{j}^{*}$ commutes with colimits and thus also the left hand side of the equivalence. Analogously, one can show that $f_{i}^{*}\circ p_{i}^{Y}$ commutes with colimits. Therefore, it is enough to show that
@@
The left hand side is equivalence to $\mathop{\textup{colim}}_{i\rightarrow k,j\rightarrow k}X_{ik*}\circ X_{jk}^{*}\circ f_{j}^{*}$ by the above and analogously, we see that the right hand side is equivalent to $\mathop{\textup{colim}}_{i\rightarrow k,j\rightarrow k}f_{i}^{*}\circ Y_{ik*}\circ Y_{jk}^{*}$. Since each square in $F(f)$ is right adjointable these colimits are in fact equivalent, proving the claim and therefore assertion. ∎
@@
Consider the setting of Lemma 2.20 (2). Let $f$ be a morphism in $\operatorname{Fun}(K,{\mathcal{E}})$ such that $f_{k}$ admits a left adjoint for all $k\in K$. Let $f_{\sharp}$ denote the left adjoint of $f$ given by Lemma 2.22. Then the following square
${F(X_{k})}$${F_{K}(X)}$${F(Y_{k})}$${F_{K}(Y)}$$\scriptstyle{\textup{ins}_{X}^{k}}$$\scriptstyle{f_{k\sharp}}$$\scriptstyle{f_{\sharp}}$$\scriptstyle{\textup{ins}_{Y}^{k}}$
commutes (up to homotopy) for all $k\in K$.
@@
${F_{K}(X)}$${\lim_{K}F_{*}\circ X}$${F_{K}(Y)}$${\lim_{K}F_{*}\circ Y}$$\scriptstyle{h_{X}}$$\scriptstyle{f_{\sharp}}$$\scriptstyle{f_{\sharp}}$$\scriptstyle{h_{Y}}$
@@
${\lim_{K}F_{*}\circ Y}$${F(Y_{i})}$${\lim_{K}F_{*}\circ X}$${F(X_{k})}$$\scriptstyle{p_{k}^{Y}}$$\scriptstyle{f^{*}}$$\scriptstyle{f_{k}^{*}}$$\scriptstyle{p_{k}^{X}}$
commutes (up to homotopy) for all $k\in K$. So in particular, we conclude by passing to left adjoints. ∎
@@
We are left to show (2) and (3). Invoking Lemma 2.20 resp. Lemma 2.23 and the fact that each compact of ${\mathcal{M}}(X)$ resp. ${\mathcal{M}}(Y)$ comes from some level $i\in K$ (cf. Lemma 2.19), we can reduce the questions to questions on ${\mathcal{C}}$. But on ${\mathcal{C}}$ this follows from definition. ∎
@@
Let us consider the setting of Example 2.15 (2). By Proposition 2.18 we immediately see that the functor $\textup{DM}\colon(\textup{ProAlg}_{S}^{{\mathcal{I}}})^{\operatorname{op}}\rightarrow\textup{CAlg}(\Pr^{L}_{\omega})$ defines a pullback formalism with respect to the class of adjointable smooth morphisms.
2.2 Coefficient systems
In this short subsection, we want to show that our extended pullback formalisms allow us to define coefficient systems on $-{\mathcal{I}}$-algebraic stacks. We have to be a bit more precise, when formulating homotopy invariance $T$-stability and localization. For example $\mathbb{A}^{1}$-invariance makes only sense for cartesian morphisms, i.e. the $-{\mathcal{I}}$-algebraic version of $\mathbb{A}^{1}$ needs to be pulled back from $\mathbb{A}^{1}_{S}$. This will be made more precise in the Corollary 2.26.
We first start with the definition of a coefficient system over an arbitrary scheme $Q$ following Drew (cf. [Dre18]). We want to remark that we do only work with compactly generated. The main reason is that we have to use Proposition 2.18 that we only know in this case.
Definition 2.25 ([Dre18]).
A coefficient system is a compactly generated pullback formalism
${\mathcal{M}}^{*,\otimes}\colon(\textup{AffSch}^{\kappa}_{Q})^{\operatorname{op}}\rightarrow\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega}),\quad f\mapsto f^{*}\quad$
satisfying the following conditions.
  1. (Loc)
    For any closed immersion $i\colon Z\hookrightarrow X$ in $\textup{Sch}_{Q}$ with open complement $j\colon U\hookrightarrow X$, the square
@@ 2.2 Coefficient systems
    is a pullback in $\Pr^{L}$.
  2. (Ho)
    For each $X\in\textup{Sch}_{Q}$ the *-pullback along the natural projection $p\colon\mathbb{A}^{1}_{X}\rightarrow X$ is fully faithful, i.e. the unit $\operatorname{id}\rightarrow p_{*}p^{*}$ is an equivalence.
  3. (St)
    For any smooth morphism $f\colon X\rightarrow Y$ in $\textup{Sch}_{Q}$ admitting a section $s$, the transformation $f_{\sharp}s_{*}$ is an equivalence.
Let ${\mathcal{M}}^{*\otimes}$ be a coefficient system. Then we define the class ${\mathcal{P}}_{{\mathcal{I}}}$ similarly to the definition in Proposition 2.18. To be more precise, ${\mathcal{P}}_{{\mathcal{I}}}$ consists of smooth adjointable morphisms $f$ in $\textup{ProAlg}^{{\mathcal{I}}}$. By Proposition 2.18 it is easy to see that ${\mathcal{M}}_{{\mathcal{I}}}$ defines a coefficient system with respect to the class ${\mathcal{P}}_{{\mathcal{I}}}$ in the following sense.
Corollary 2.26.
Let ${\mathcal{M}}^{*,\otimes}$ be a coefficient system that satisfies étale descent. Then ${\mathcal{M}}_{{\mathcal{I}}}$, that is the right Kan extension of ${\mathcal{M}}^{*,\otimes}$ along Yoneda to Artin $S$-stacks locally of finite type and then the composition of $\operatorname{Fun}(K,{\mathcal{M}}^{*\otimes})$ with the colimit functor, is a pullback formalism in the following sense.
  1. (PF)
    The functor ${\mathcal{M}}_{{\mathcal{I}}}$ defines a pullback formalism for the class ${\mathcal{P}}_{{\mathcal{I}}}$.
  2. (Loc)
    Let $i\colon(z,Z)\hookrightarrow(x,X)$ be cartesian closed immersion of $-{\mathcal{I}}$-algebraic stacks with open complement $j\colon(u,U)\rightarrow(x,X)$. Then there following diagram
@@
    is a pullback in $\Pr^{L}$.
  3. (Ho)
    Let $(x,X)$ be a $-{\mathcal{I}}$-algebraic stack and let us define $(q,\mathbb{A}^{1}_{X})$ via the system $n\mapsto\mathbb{A}^{1}_{X_{i}}$. Then $p^{*}$ is fully faithful, where $p\colon(q,\mathbb{A}^{1}_{X})\rightarrow(x,X)$ denotes the projection.
  4. (St)
    For any smooth adjointable morphism $f\colon(x,X)\rightarrow(y,Y)$ of $-{\mathcal{I}}$-algebraic stacks, admitting a section $s$, the transformation $f_{\sharp}s_{*}$ is an equivalence.
Proof.
Claim (PF) follows from Proposition 2.18. Similar to the proof of (2) and (3) in Proposition 2.18, we can reduce to the case of Artin stacks via Lemma 2.20 and Lemma 2.19. For Artin stacks this follows immediately from descent. ∎
Remark 2.27.
The proof of Corollary 2.26 shows that if we restrict ourselves to pro-schemes locally of finite type over $S$, then we may drop the descent assumption on our coefficient system.
2.3 Six functor formalisms
In this section, we want to prove Theorem 5. For this, we need to extend $6$-functor formalisms to pro-algebraic stacks. We will freely use the formalism of Gaitsgory-Rozenblym, Liu-Zheng and Mann (cf. [GR17], [LZ12], [Man22]). We will use the results of the latter and for details refer to [Man22, §A.5]. We do not want to recall the main definitions and freely use the language of geometric setups and correspondences as in loc. cit..
Let $({\mathcal{C}},E)$ be a geometric setup, that is an $\infty$-category ${\mathcal{C}}$ and a class of morphisms $E$ in ${\mathcal{C}}$ that is closed under pullbacks, compositions and equivalences, such that ${\mathcal{C}}$ admits finite limits. Let $K$ a filtered $\infty$-category. Let ${\mathcal{D}}\colon\textup{Corr}({{\mathcal{C}}})_{E}\rightarrow\textup{Cat}_{\infty}$ be a $6$-functor formalism in the sense of [Man22, §A.5], where $\textup{Corr}({{\mathcal{C}}})_{E}$ is the $\infty$-category of correspondences in the sense of [LZ12].
We let ${\mathcal{D}}_{{\mathcal{C}}}\colon{\mathcal{C}}^{\operatorname{op}}\rightarrow\textup{CAlg}(\textup{Cat}_{\infty})$ denote the restriction of ${\mathcal{D}}$ under the inclusion ${\mathcal{C}}^{\operatorname{op}}\hookrightarrow\textup{Corr}({{\mathcal{C}}})_{E}$. We assume that ${\mathcal{D}}_{{\mathcal{C}}}$ factors through $\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega})$. We will use a similar notation as in Section 2.1. We denote by ${\mathcal{D}}_{K}$ the composition
[…]
We define $E_{K}$ to be the class of those morphisms $f\colon X\rightarrow Y$ in $\operatorname{Fun}(K,{\mathcal{C}}^{\operatorname{op}})$ such that $f_{k}$ is in $E$ for all $k\in K$ and for all $k\rightarrow l\in K$ the square
[…]
@@ 2.3 Six functor formalisms
Assume that $K$ admits an initial object 00 and $E$ admits a suitable decomposition $J,F$ that satisfies the assumptions of [Man22, Prop. A.5.10]. Then ${\mathcal{D}}_{K}$ can be upgraded to a $6$-functor formalism
${\mathcal{D}}_{K}\colon\textup{Corr}({\operatorname{Fun}(K^{\operatorname{op}},{\mathcal{C}})})_{E_{K}}\rightarrow\textup{CAlg}(\textup{Pr}^{L,\otimes}_{\omega}).$
Further, if $f\in E$ is a morphism such that for all $k\in K$ we have $f_{k}\in F$ resp. $f(k)\in J$, then $f_{!}\simeq f_{*}$ resp. $f_{!}$ is left adjoint to $f^{*}$.
@@
We define $J_{K}$ resp. $F_{K}$ to be the class of those morphisms $f$ in $E_{K}$ such that $f(k)$ is in $J$ resp. $F$ for all $k\in K$. Then we claim that $J_{K},F_{K}$ is a suitable decomposition of $E$.
Indeed, since for a map $f$ in $E$ all squares of the associated diagram are pullback squares. Thus, all of the properties follow immediately (note that we need that $K$ admits an initial object for the decomposition property).
@@
By Proposition 2.18 the functor ${\mathcal{D}}_{K}$ defines a pullback formalism on $\operatorname{Fun}(K,{\mathcal{C}})$ with respect to $J_{K}$.
By construction, the pullback $f^{*}$ along any morphism $f\in\operatorname{Fun}(K,{\mathcal{C}})$ admits a right adjoint. We need to check that for morphisms in $F_{K}$ we have base change and projection formula. But similar to the proof of Proposition 2.18, we can use that any compact object comes from some compact on the $k$-th level, for some $k\in K$ (by mutas mutandis of Lemma 2.19). This argument shows that for any pullback square
${W}$${X}$${Y}$${Z}$$\scriptstyle{f^{\prime}}$$\scriptstyle{g^{\prime}}$$\scriptstyle{g}$$\scriptstyle{f}$
with $g\in J_{K}$ and $f\in F_{K}$ we have $g_{\sharp}f^{\prime}_{*}\simeq f_{*}g^{\prime}_{\sharp}$ completing the proof. ∎
The above proposition a priori only applies if the geometric setup has a suitable decomposition. For schemes this is usually not a problem as in practice we can decompose separated morphisms via compactifications into open immersions and proper maps. For algebraic stacks this only holds locally. Nevertheless, we can extend the functors by the base change formula to cartesian morphisms of $-{\mathcal{I}}$-algebraic stacks (cf. Proposition 2.29 below).
@@
Let $f\colon X\rightarrow Y$ be in $E_{K}$. Then there exists an adjunction
${f_{!}\colon{\mathcal{D}}_{K}(X)}$${{\mathcal{D}}_{K}(Y)\colon f^{!}}$
satisfying base change and projection formula with respect to $f^{*}\coloneqq{\mathcal{D}}_{K}(f)$.

### v4

@@ 3.1 Pullback formalisms
$F(Y_{j})$$F_{{\mathcal{I}}}(Y)$$\lim_{{\mathcal{I}}}F_{*}\circ Y$$F(X_{j})$$F_{{\mathcal{I}}}(X)$$\lim_{{\mathcal{I}}}F_{*}\circ X$$\scriptstyle{\textup{ins}_{Y}^{j}}$$\scriptstyle{f_{j}^{*}}$$\scriptstyle{f^{*}}$$\scriptstyle{h_{Y}}$$\scriptstyle{f^{*}}$$\scriptstyle{\textup{ins}_{X}^{j}}$$\scriptstyle{h_{X}}$
for any $j\in{\mathcal{I}}$. First, we show that $p_{i}^{Y}\circ f^{*}$ commutes with colimits. Indeed, it is enough to see that $p_{i}^{Y}\circ f^{*}\circ h_{X}\circ\textup{ins}_{X}^{j}$ commutes with colimits, then is equivalent to the induced map $F_{{\mathcal{I}}}(X)\rightarrow F(Y_{i})$ in $\Pr^{L}$, which necessarily commutes with colimits. By Lemma 2.10 (2), we have
$p_{i}^{X}\circ f^{*}\circ h_{Y}\circ\textup{ins}_{Y}^{j}\simeq p_{i}^{Y}\circ h_{X}\circ\textup{ins}_{X}^{j}\circ f_{j}^{*}\simeq\mathop{\textup{colim}}_{i\rightarrow k,j\rightarrow k}X_{ik*}\circ X_{jk}^{*}\circ f_{j}^{*}$
(recall our notations in the introduction under Section 1.1). By construction each of the functors $X_{ik*},X_{jk}^{*},f_{j}^{*}$ commutes with colimits and thus also the left hand side of the equivalence. Analogously, one can show that $f_{i}^{*}\circ p_{i}^{Y}$ commutes with colimits. Therefore, it is enough to show that
@@
The left hand side is equivalent to $\mathop{\textup{colim}}_{i\rightarrow k,j\rightarrow k}X_{ik*}\circ X_{jk}^{*}\circ f_{j}^{*}$ by the above and analogously, we see that the right hand side is equivalent to $\mathop{\textup{colim}}_{i\rightarrow k,j\rightarrow k}f_{i}^{*}\circ Y_{ik*}\circ Y_{jk}^{*}$. Since each square in $F(f)$ is right adjointable these colimits are in fact equivalent, proving the claim and therefore assertion. ∎
@@
Let $f$ be a morphism in $\operatorname{Fun}({\mathcal{I}},{\mathcal{E}}^{\operatorname{op}})$ such that $f_{i}$ admits a left adjoint for all $i\in{\mathcal{I}}$. Let $f_{\sharp}$ denote the left adjoint of $f$ given by Lemma 3.9. Then the following square
$F(X_{i})$$F_{{\mathcal{I}}}(X)$$F(Y_{i})$$F_{{\mathcal{I}}}(Y)$$\scriptstyle{\textup{ins}_{X}^{i}}$$\scriptstyle{f_{i\sharp}}$$\scriptstyle{f_{\sharp}}$$\scriptstyle{\textup{ins}_{Y}^{i}}$
commutes (up to homotopy) for all $i\in{\mathcal{I}}$.
@@
$F_{{\mathcal{I}}}(X)$$\lim_{{\mathcal{I}}}F_{*}\circ X$$F_{{\mathcal{I}}}(Y)$$\lim_{{\mathcal{I}}}F_{*}\circ Y$$\scriptstyle{h_{X}}$$\scriptstyle{f_{\sharp}}$$\scriptstyle{f_{\sharp}}$$\scriptstyle{h_{Y}}$
@@
$\lim_{{\mathcal{I}}}F_{*}\circ Y$$F(Y_{i})$$\lim_{{\mathcal{I}}}F_{*}\circ X$$F(X_{i})$$\scriptstyle{p_{i}^{Y}}$$\scriptstyle{f^{*}}$$\scriptstyle{f_{i}^{*}}$$\scriptstyle{p_{i}^{X}}$
commutes (up to homotopy) for all $i\in{\mathcal{I}}$. So in particular, we conclude by passing to left adjoints. ∎
@@
We are left to show (2) and (3). Invoking Lemma 2.10 resp. Lemma 3.10 and the fact that each compact of ${\mathcal{M}}_{{\mathcal{I}}}(X)$ comes from some level $i\in{\mathcal{I}}$ (cf. Lemma 3.8). So, we can check (2) and (3) on some $k\in{\mathcal{I}}$ (note that we use that $\mathcal{I}$} is filtered) and this follows from the definition of $\mathcal{M}$$^{\otimes}$. ∎
3.2 Six functor formalism
In this section, we want to prove Theorem 2.9. We will first give a proof in the $(\dagger)$ case. Afterwards, we will give a proof in the $(\ast)$ case and complete the proof.
Before we continue, we want to show that properties (M1)-(M5) are a formal consequence of the compact generation of $\textup{DM}(x,X)$ (see Lemma 3.8).
Lemma 3.11.
In the setting of Theorem 2.9, assuming $f_{!}\dashv f^{!}$ exists, satisfying (F1)-(F4), the properties (M1) and (M2) hold. If $\mathcal{I}$} admits an initial object $0\in{\mathcal{I}}$, then also (M3) - (M5) hold.
Proof.
Arguing as in the proof of Proposition 3.7 and using descent the properties (M1) - (M3) reduce to questions of DM on schemes, which is well known [CD19]. The property (M4) follows from the localization sequence.
If $f$ is cartesian, smooth and then the relative dimension of $f_{i}$ and $f_{0}$ agree for all $i\in{\mathcal{I}}$. From here (M5) follows similarly as before. ∎
3.2.1 Proof of Theorem 2.9 for adjointable maps
In this subsection, we will assume that we are in the setting $(\dagger)$ and we simply say pro-algebraic instead of pro-$\mathbb{N}_{0}$-algebraic.
Lemma 3.12.
Let $\mathcal{C}$} be an $\infty$-category. Let $\textup{Spine}[\mathbb{N}_{0}]$ be the simplicial subset of $\mathbb{N}_{0}$ of finite vertices that are joined by edges (or more informally, we forget about all $n\geq 2$ simplices in $\mathbb{N}_{0}$). Then the restriction $\operatorname{Fun}(\mathbb{N}_{0},{\mathcal{C}})\rightarrow\operatorname{Fun}(\textup{Spine}[\mathbb{N}_{0}],{\mathcal{C}})$ is a trivial Kan fibration.
In particular, any natural transformation of functors $\alpha\colon\textup{Spine}[\mathbb{N}_{0}]\times\Delta^{1}\rightarrow{\mathcal{C}}$ can be lifted to a natural transformation in $\hat{\alpha}\colon\mathbb{N}_{0}\times\Delta^{1}\rightarrow{\mathcal{C}}$.
Proof.
Indeed, the functor $\textup{Spine}[\mathbb{N}_{0}]\rightarrow\mathbb{N}_{0}$ is inner anodyne [Lur21, 03HK] proving the claim. ∎
Proposition 3.13.
Let $(x,X)$ be a tame pro-algebraic stack, then
$\textup{DM}(x,X)\simeq\mathop{\textup{colim}}_{i\in\mathbb{N}_{0},x^{!}}\textup{DM}^{!}(X_{i}).$
Proof.
By descent, we see that for each $n\in\mathbb{N}_{0}$ we have $\tilde{p}_{i}\colon\textup{DM}^{!}(X_{i})\simeq\textup{DM}^{*}(X_{i})$ as $X_{i}$ admits a smooth covering by a scheme. We will construct equivalences
$p_{i}\colon\textup{DM}^{!}(X_{i})\simeq\textup{DM}^{*}(X_{i})$
out of $\tilde{p}_{n}$ that are compatible with the presentation as a colimit. This is easily achieved by twisting with the Thom-motives associated to the smooth maps $x_{0i}$. But let us be more precise.
Let us define $p_{i}\coloneqq\tilde{p}_{n}\otimes\textup{Th}(\Omega_{x_{0n}})1_{X_{n}}$, where Th denotes the Thom-motive [^2] Let $p\colon V\rightarrow X_{i}$ be a vector bundle with zero section $s$, then $\textup{Th}(V)\coloneqq p_{\sharp}s_{*}$. associated to $\mathbb{V}(\Omega_{x_{0i}})\rightarrow X_{n}$. Then $p_{n}$ is still an equivalence of $\textup{DM}^{!}(X_{n})$ and $\textup{DM}^{*}(X_{n})$. By compatibility with the purity equivalence with composition [CD19, Rem. 2.4.52], the $p_{i}$ are compatible with the presentation of $\textup{DM}(x,X)$ as a colimit by Lemma 3.12. ∎
@@ 3.2 Six functor formalism
To generalize Proposition 3.13 to arbitrary $\mathcal{I}$} it is enough to know that the level-wise equivalences $p_{i}$ in the proof yield an equivalence of diagrams. While on the homotopy categorical level this is clear, we were not able to write a map down for higher simplices.
Proof of Theorem 2.9.
The property (PB) is Proposition 3.7. By Proposition 3.13, we obtain a functor $f^{!}\colon\textup{DM}(y,Y)\rightarrow\textup{DM}(x,X)$. Applying Proposition 3.7 to $f^{!}$, we obtain a left adjoint $f_{!}$ of $f^{!}$. Note that we use that each $f_{n}^{!}$ preserves colimits, since $f_{n!}$ preserves compact objects. In particular, (F2) follows from construction. The properties (F3) and (F4) can be checked on compacts, which is clear by design and Lemma 3.8.
The rest follows from Lemma 3.11
Remark 3.15.
Let $(x,X)$ be a strict tame pro-algebraic stack and $(y,Y)$ a tame pro-algebraic stack together with a morphism $f\colon(x,X)\to(y,Y)$. Assume that for $y_{n}\colon Y_{n+1}\to Y_{n}$ and $n\geq 0$ the unit $\operatorname{id}\to y_{n}^{*}y_{n\sharp}$ is an equivalence. In this setting, we can make $f_{!}$ more explicit.
For each $n\geq 0$, we obtain a functor $f_{n!}$. We want to construct a morphism of diagrams $f_{\bullet!}\colon\mathbb{N}_{0}\times\Delta^{1,\operatorname{op}}\rightarrow\textup{DGCat}_{\textup{cont}}$, which after taking the colimit then yields $f_{!}$. Since our morphisms are not cartesian, we have to take the Thom-twist of $x_{n}\colon X_{n+1}\to X_{n}$ and $y_{n}$ into account. Set $\tilde{f}_{n!}\coloneqq(\mathrm{Th}(\Omega_{y_{0n}})1_{Y_{n}})^{-1}\otimes f_{n!}\circ\mathrm{Th}(\Omega_{x_{0n}})1_{X_{n}}\otimes$-. Then, we obtain
$y_{n}^{*}\tilde{f}_{n!}$ $\simeq y_{n}^{*}(\mathrm{Th}(\Omega_{y_{0n}})1_{Y_{n}})^{-1}\otimes f_{n!}\circ\mathrm{Th}(\Omega_{x_{0n}})1_{X_{n}}\otimes x_{n!}x_{n}^{!}$
$\simeq y_{n}^{*}(\mathrm{Th}(\Omega_{y_{0n}})1_{Y_{n}})^{-1}\otimes y_{n!}f_{n+1!}\circ\mathrm{Th}(\Omega_{x_{0n+1}})1_{X_{n+1}}\otimes x_{n}^{*}$
$\simeq y_{n}^{*}y_{n\sharp}\circ(\mathrm{Th}(\Omega_{y_{0n+1}})1_{Y_{n+1}})^{-1}\otimes f_{n+1!}\circ\mathrm{Th}(\Omega_{x_{0n+1}})1_{X_{n+1}}\otimes x_{n}^{*}\simeq\tilde{f}_{n+1!}x_{n}^{*},$
where we use compatibility of Thom twists with composition in the sense of [CD19, Rem. 2.4.52] and the projection formula.
In particular, we obtain the desired diagram $f_{\bullet!}\colon\mathbb{N}_{0}\times\Delta^{1,\operatorname{op}}\rightarrow\textup{DGCat}_{\textup{cont}}$ that after taking colimits induces a functor $f_{!}\colon\textup{DM}(x,X)\rightarrow\textup{DM}(y,Y)$. This is a colimit preserving functor and therefore admits a right adjoint. By construction of the comparison morphism in Proposition 3.13, this agrees with the construction of $f_{!}$ in the proof of Theorem 2.9.
This construction also shows why (F1) fails in this setting. Even when $f$ is proper, the resulting $f_{!}$ is an infinite twist of $f_{*}$. Note however, if $f$ is cartesian, then by design the twists appearing in $\tilde{f}$ cancel each other out, so this caveat appears only in the non-cartesian setting.
3.2.2 Proof of Theorem 2.9 for cartesian maps
We now proceed to prove Theorem 2.9 in the $(\ast)$ case. Throughout, we use the term pro-algebraic in place of pro-$\mathcal{I}$}-algebraic.
To that end, we need to extend the six-functor formalism to pro-algebraic stacks. We will freely use the frameworks developed by Gaitsgory–Rozenblyum, Liu–Zheng, and Mann (cf. [GR17], [LZ17], [Man22]). Our arguments rely on results from the latter, and for details we refer the reader to [Man22, §A.5]. We will not recall the main definitions here, and instead use the language of geometric setups and correspondences as developed in loc. cit.
Let $({\mathcal{C}},E)$ be a geometric setup, meaning an $\infty$-category $\mathcal{C}$} admitting finite limits, together with a class of morphisms $E\subset\text{Mor}({\mathcal{C}})$ that is stable under pullbacks, compositions, and equivalences. Let
${\mathcal{D}}\colon\textup{Corr}({{\mathcal{C}}})_{E}\to\textup{Cat}_{\infty}$
be a six-functor formalism in the sense of [Man22, §A.5], where $\textup{Corr}({{\mathcal{C}}})_{E}$ denotes the $\infty$-category of correspondences as defined in [LZ17].
We denote by
${\mathcal{D}}_{{\mathcal{C}}}\colon{\mathcal{C}}^{\operatorname{op}}\to\textup{CAlg}(\textup{Cat}_{\infty})$
@@
We denote by $\mathcal{D}$$_{{\mathcal{I}}}$ the composition
@@
We define $E_{{\mathcal{I}}}$ to be the class of those morphisms $f\colon X\rightarrow Y$ in $\operatorname{Fun}({\mathcal{I}},{\mathcal{C}}^{\operatorname{op}})$ such that $f_{i}$ is in $E$ for all $i\in{\mathcal{I}}$ and for all $i\rightarrow j\in{\mathcal{I}}$ the square
$X_{j}$$X_{i}$$Y_{j}$$Y_{i}$$\scriptstyle{X_{ij}}$$\scriptstyle{f_{j}}$$\scriptstyle{f_{i}}$$\scriptstyle{Y_{ij}}$
@@
Assume that $\mathcal{I}$} admits an initial object 0 and $E$ admits a suitable decomposition $J,F$ that satisfies the assumptions of [Man22, Prop. A.5.10]. Then $\mathcal{D}$$_{{\mathcal{I}}}$ can be upgraded to a $6$-functor formalism
[…]
Further, if $f\in E$ is a morphism such that for all $i\in{\mathcal{I}}$ we have $f_{i}\in F$ resp. $f(i)\in J$, then $f_{!}\simeq f_{*}$ resp. $f_{!}$ is left adjoint to $f^{*}$.
@@
We define $J_{{\mathcal{I}}}$ resp. $F_{{\mathcal{I}}}$ to be the class of those morphisms $f$ in $E_{{\mathcal{I}}}$ such that $f(i)$ is in $J$ resp. $F$ for all $i\in{\mathcal{I}}$. Then we claim that $J_{{\mathcal{I}}},F_{{\mathcal{I}}}$ is a suitable decomposition of $E$.
Indeed, since for a map $f$ in $E$ all squares of the associated diagram are pullback squares. Thus, all of the properties follow immediately (note that we need that $\mathcal{I}$} admits an initial object for the decomposition property).
@@
By Proposition 3.7 the functor $\mathcal{D}$$_{{\mathcal{I}}}$ defines a pullback formalism on $\operatorname{Fun}({\mathcal{I}},{\mathcal{C}}^{\operatorname{op}})$ with respect to $J_{{\mathcal{I}}}$.
By construction, the pullback $f^{*}$ along any morphism $f\in\operatorname{Fun}({\mathcal{I}},{\mathcal{C}}^{\operatorname{op}})$ admits a right adjoint. We need to check that for morphisms in $F_{{\mathcal{I}}}$ we have base change and projection formula. But similar to the proof of Proposition 3.7, we can use that any compact object comes from some compact on the $i$-th level, for some $i\in{\mathcal{I}}$ (by mutas mutandis of Lemma 3.8). This argument shows that for any pullback square
[…]
with $g\in J_{{\mathcal{I}}}$ and $f\in F_{{\mathcal{I}}}$ we have $g_{\sharp}f^{\prime}_{*}\simeq f_{*}g^{\prime}_{\sharp}$ completing the proof. ∎
The above proposition a priori only applies if the geometric setup has a suitable decomposition. For schemes this is usually not a problem as in practice we can decompose separated morphisms via compactifications into open immersions and proper maps. For algebraic stacks this only holds locally. Nevertheless, the existence of the !!-adjunction extends to cartesian morphisms of pro-$\mathcal{I}$}-algebraic stacks by the base change formula, as seen below.
@@
Let $f\colon X\rightarrow Y$ be a morphism inside $E_{{\mathcal{I}}}\subseteq\operatorname{Fun}({\mathcal{I}}^{\operatorname{op}},{\mathcal{C}})^{\operatorname{op}}$. Then there exists an adjunction
$f_{!}\colon{\mathcal{D}}_{{\mathcal{I}}}(X)$${\mathcal{D}}_{{\mathcal{I}}}(Y)\colon f^{!}$}
satisfying base change and projection formula with respect to $f^{*}\coloneqq{\mathcal{D}}_{{\mathcal{I}}}(f)$.